# Pseudo-real quantum fields

### v1

@@ 3.5 Finding $\eta$
This equation is non-linear, but it can be solved perturbatively in orders of the coupling constant $g$. Thus, at order $g^{2}$, the map $\eta_{P}$ is of the form
$\eta_{P}[\phi(x)]=-\phi(x)+ig\int{\rm d}^{4}y\>G_{0}(x,y)\phi^{2}(y)+g^{2}\int{\rm d}^{4}y\>{\rm d}^{4}z\>G_{0}(x,y)\phi(y)G_{0}(y,z)\phi^{2}(z)+\mathcal{O}(g^{3})\>.$ (3.29)
@@
Here, the two-point function $G_{0}(x,y)$ is the Green’s function (3.11) of the free Klein-Gordon equation. Note, however, that the field $\phi$ is not the free Klein-Gordon field $\phi_{0}$, but the complex interacting field of our Lagrangian, which itself can be expanded in orders of $g$ [see Eq. (3.10)]. The map $\eta_{P}$ bears some resemblance to the map derived in Refs. [22, 23]. [^1] In Ref. [22], it is suggested that there exists an arbitrariness to the Hermitian theory that corresponds to the $i\phi^{3}$ theory in its $PT$-symmetric regime. We disagree with this conclusion, and it is readily confirmed that the apparent arbitrariness remaining in Eq. (17) of Ref. [22] is eliminated when performing the path integral. This is to say that the terms proportional to the unconstrained coefficients $c_{2}$, $c_{4}$ and $c_{5}$ that appear therein cancel at order $g^{2}$.
@@
The pseudo-reality condition (3.27) gives the relationship between the field $\phi$ and its Hermitian conjugate $\phi^{\dagger}$:
$\phi^{{\dagger}}(x)=\phi(x)-ig\int{\rm d}^{4}y\>G_{0}(x,y)\phi^{2}(y)-g^{2}\int{\rm d}^{4}y\>{\rm d}^{4}z\>G_{0}(x,y)\phi(y)G_{0}(y,z)\phi^{2}(z)+\mathcal{O}(g^{3})$ (3.30)
@@
This equation defines a hypersurface in the complex field space and reduces the number of degrees of freedom of the complex field $\phi$ to that of a real field.
@@
The time-reversal map $\eta_{T}$ solves the equation of motion for $-\eta_{P}$. Hence, the solution is straightforward
$\eta_{T}[\phi(x)]=\phi(x)-ig\int{\rm d}^{4}y\>G_{0}(x,y)\phi^{2}(y)-g^{2}\int{\rm d}^{4}y\>{\rm d}^{4}z\>G_{0}(x,y)\phi(y)G_{0}(y,z)\phi^{2}(z)\>+\mathcal{O}(g^{2})\;,$ (3.31)
and this yields the same pseudo-reality condition as Eq. (3.30).
3.6 $\eta$ operator
@@
$\eta_{P}[\hat{\phi}(x)]=\hat{\eta}_{P}\hat{\phi}(x_{P})\hat{\eta}^{-1}_{P}\qquad\text{and}\qquad\eta_{T}[\hat{\phi}(x)]=\hat{\eta}_{T}\hat{\phi}(x_{T})\hat{\eta}^{-1}_{T}\>.$ (3.32)
First, we need to choose a Green’s function $G_{0}(x,y)$ in the expansion (3.29). The equal-time commutation relations
$[\hat{\phi}(t,\vec{x}),\hat{\pi}(t,\vec{y})]=i\delta^{(3)}(\vec{x}-\vec{y})\>,$ (3.33a)
$[\hat{\phi}(t,\vec{x}),\hat{\phi}(t,\vec{y})]=[\hat{\pi}(t,\vec{x}),\hat{\pi}(t,\vec{y})]=0\>,$ (3.33b)
are only preserved for a Hermitian Green’s function. Here, the operator $\pi(x)$ is the conjugate momentum $\pi(x)=\frac{\partial\mathscr{L}(x)}{\partial\dot{\phi}(x)}=\dot{\phi}(x)$ of the interacting field $\phi$. However, we note that the field operator $\hat{\phi}(x)$ and its Hermitian conjugate $\hat{\phi}^{\dagger}(x)$ do not commute at equal times:
$[\hat{\phi}(t,\vec{x}),\hat{\phi}^{\dagger}(t,\vec{y})]\neq 0\>.$ (3.34)
It should not be surprising as $\phi^{\dagger}$ is not the dual field to $\phi$. Indeed, the dual field $\tilde{\phi}^{\dagger}$ commutes with $\phi$ at equal times:
$[\hat{\phi}(t,\vec{x}),\hat{\tilde{\phi}}^{{\dagger}}(t,\vec{y})]=0\>.$ (3.35)
The only Hermitian Green’s functions solving (3.11) are the retarded and the advanced propagators:
[…]
@@
$D(x,y)=\int\!\frac{{\rm d}^{3}\vec{p}}{(2\pi)^{3}}\>\frac{1}{2iE_{\vec{p}}}\,e^{-ip\cdot(x-y)}\>.$ (3.37)
@@
We can show that the operator $\hat{\eta}_{P}=\hat{\eta}_{g}\hat{\mathcal{P}}$ can be written as a combination of the operator $\hat{\eta}_{g}$, yielding the series expansion in $g$, and the ‘standard’ parity operator $\hat{\mathcal{P}}$. The ‘standard’ parity operator has the usual form
$\hat{\mathcal{P}}(t)\hat{\phi}(t,\vec{x})\hat{\mathcal{P}}^{-1}(t)=-\hat{\phi}(t,-\vec{x})\>,\qquad\hat{\mathcal{P}}(t)=e^{\frac{\pi}{2}\int d^{3}z(\hat{\phi}(t,\vec{z})\hat{\pi}(t,-\vec{z})+\hat{\phi}(t,\vec{z})\hat{\pi}(t,\vec{z}))}\>.$ (3.38)
We note that the Lagrangian is pseudo-Hermitian with respect to $\hat{\eta}_{P}$, but not with respect to the $\hat{\mathcal{P}}$ operator.
The operator $\hat{\eta}_{g}$ is less trivial, and the full calculation is provided in Appendix A. We are looking for an expansion of the form
@@
$+ig^{2}\int^{t}_{t_{0}}{\rm d}y^{0}\int^{y^{0}}_{t_{0}}{\rm d}z^{0}\int{\rm d}^{3}\vec{y}\>{\rm d}^{3}\vec{z}\>G_{R}(x,y)G_{R}(y,z)G_{R}(y,z)\hat{\phi}(z)\>.$ (3.39)
The last term is a quantum loop correction of order $\hbar$. As we take the classical limit $\hbar\mapsto 0$, this term vanishes and we recover the classical expansion in Eq. (3.29).
Given that the solution is similar to that of the interacting Klein-Gordon equation, we find that the series expansion is generated by an operator similar to the time-evolution operator; specifically,
$\hat{\eta}_{g}(t,t_{0})=\hat{U}^{-1}\quantity(-\frac{1}{3},t,t_{0})\>,$ (3.40)
@@
$\hat{U}\quantity(\alpha,t,t_{0})\coloneqq T\quantity[e^{\alpha g\int^{t}_{t_{0}}{\rm d}y^{0}\int{\rm d}^{3}\vec{y}\>\hat{\phi}^{3}(y^{0},\vec{y})}]\text{\>\>\>and\>\>\>}\hat{U}^{-1}\quantity(\alpha,t,t_{0})=\bar{T}\quantity[e^{-\alpha g\int^{t}_{t_{0}}{\rm d}y^{0}\int{\rm d}^{3}\vec{y}\>\hat{\phi}^{3}(y^{0},\vec{y})}]\>,$ (3.41)
for some constant $\alpha\in\mathbb{C}$. Here, $T$ is the time-ordering operator and $\bar{T}$ is the anti-time ordering operator. Note that $\hat{U}(\frac{1}{3!},t,t_{0})$ is (when replacing $\hat{\phi}$ by $\hat{\phi}_{0}$) the usual time-evolution operator whose action on the free Klein-Gordon field $\phi_{0}$ gives the series expansion in Eq. (3.10).
@@
$\hat{\mathcal{P}}^{\dagger}(t)=\hat{\eta}_{P}(t,t_{0})\hat{\mathcal{P}}(t)\hat{\eta}_{P}^{-1}(t,t_{0})\;.$ (3.43)
@@
and the operator $\hat{\eta}_{g}$ in Eq. (3.40). Thus, the Lagrangian is anti-pseudo-Hermitian with respect to the anti-linear operator $\hat{\eta}_{T}$:
$\hat{\eta}_{T}(t,t_{0})=\hat{\eta}_{g}(t,t_{0})\hat{\mathcal{T}}(t).$ (3.46)
@@
Given the Lagrangian is $\eta_{P}^{-1}\eta_{T}$-symmetric (3.24), we see that it is also $PT$-symmetric with respect to the ‘standard’ $PT$ transformation as $\hat{\eta}_{P}^{-1}\hat{\eta}_{T}=\hat{\mathcal{P}}\hat{\mathcal{T}}$.
@@
Thus, we have demonstrated that by adopting a first-principles approach and introducing the pseudo-reality condition, we achieve a self-consistent formulation of the $i\phi^{3}$ theory. This framework resolves the contradictions arising from Euler-Lagrange equations and ‘standard’ parity and time-reversal transformations and ensures that the theory remains well-defined with a real spectrum under exact $PT$ symmetry.
4 Generalisation to higher spins

### v2

@@ III.5 Finding $\eta$
This equation is non-linear, but it can be solved perturbatively in orders of the coupling constant $g$. Expanding $\eta_{P}=\eta_{P}^{(0)}+g\eta_{P}^{(1)}+g^{2}\eta_{P}^{(2)}$ at order $g^{2}$ gives us
$\quantity(\Box_{x}+m^{2})\eta_{P}^{(0)}-\frac{ig}{2}\quantity(\eta_{P}^{(0)})^{2}=0\>,$ (46a)
$\quantity(\Box_{x}+m^{2})\eta_{P}^{(1)}+i\quantity(\eta^{(0)})^{2}_{P}=0\>,$ (46b)
$\quantity(\Box_{x}+m^{2})\eta_{P}^{(2)}+i\eta^{(0)}_{P}\eta^{(1)}_{P}=0\>,$ (46c)
where we have added and subtracted a term $ig\quantity(\eta_{P}^{(0)})^{2}$ in the overall expansion, so that the equation for $\eta_{P}^{(0)}$ is dependent on $g$. We do this because we wish to find a functional $\eta_{P}[\phi(x)]$ which depends on the complex interacting field $\phi$ and not on the real free Klein-Gordon field $\phi_{0}$.
The equation (46a) has a solution $\eta_{P}^{(0)}[\phi(x)]=-\phi(x)$. The remaining equations are solved using an Ansatz and, at order $g^{2}$, the map $\eta_{P}$ is found to be of the form
$\eta_{P}[\phi(x)]=-\phi(x)+ig\int{\rm d}^{4}y\>G_{0}(x,y)\phi^{2}(y)+g^{2}\int{\rm d}^{4}y\>{\rm d}^{4}z\>G_{0}(x,y)\phi(y)G_{0}(y,z)\phi^{2}(z)+\mathcal{O}(g^{3})\>.$ (47)
@@
Here, the two-point function $G_{0}(x,y)$ is the Green’s function of the free Klein-Gordon equation (27). The map $\eta_{P}$ bears some resemblance to the map derived in Refs. Li:2023kpi ; Li:2024xms . [^1] In Ref. Li:2023kpi , it is suggested that there exists an arbitrariness to the Hermitian theory that corresponds to the $i\phi^{3}$ theory in its $PT$-symmetric regime. We disagree with this conclusion, and it is readily confirmed that the apparent arbitrariness remaining in Eq. (17) of Ref. Li:2023kpi is eliminated when performing the path integral. This is to say that the terms proportional to the unconstrained coefficients $c_{2}$, $c_{4}$ and $c_{5}$ that appear therein cancel at order $g^{2}$.
@@
The pseudo-reality condition (44) gives the relationship between the field $\phi$ and its Hermitian conjugate $\phi^{\dagger}$:
$\phi^{{\dagger}}(x)=\phi(x)-ig\int{\rm d}^{4}y\>G_{0}(x,y)\phi^{2}(y)-g^{2}\int{\rm d}^{4}y\>{\rm d}^{4}z\>G_{0}(x,y)\phi(y)G_{0}(y,z)\phi^{2}(z)+\mathcal{O}(g^{3})\;.$ (48)
@@
This equation defines a hypersurface in the complex field space and reduces the number of degrees of freedom of the complex field $\phi$ to that of a real field, i.e., one.
@@
The time-reversal-like map $\eta_{T}$ solves the equation of motion for $\eta_{P}$. Hence, the solution is straightforward
$\eta_{T}[\phi(x)]=\phi(x)-ig\int{\rm d}^{4}y\>G_{0}(x,y)\phi^{2}(y)-g^{2}\int{\rm d}^{4}y\>{\rm d}^{4}z\>G_{0}(x,y)\phi(y)G_{0}(y,z)\phi^{2}(z)\>+\mathcal{O}(g^{2})\;,$ (49)
and yields the same pseudo-reality condition as Eq. (48).
III.6 $\hat{\eta}$ operator
@@
$\eta_{P}[\hat{\phi}(x)]=\hat{\eta}_{P}\hat{\phi}(x_{P})\hat{\eta}^{-1}_{P}\qquad\text{and}\qquad\eta_{T}[\hat{\phi}(x)]=\hat{\eta}_{T}\hat{\phi}(x_{T})\hat{\eta}^{-1}_{T}\>.$ (50)
First, we need to choose a Green’s function $G_{0}(x,y)$ in the expansion (47). The equal-time commutation relations
$[\hat{\phi}(t,\vec{x}),\hat{\pi}(t,\vec{y})]=i\delta^{(3)}(\vec{x}-\vec{y})\>,$ (51a)
$[\hat{\phi}(t,\vec{x}),\hat{\phi}(t,\vec{y})]=[\hat{\pi}(t,\vec{x}),\hat{\pi}(t,\vec{y})]=0\>,$ (51b)
are only preserved under the transformation $\eta$ for a causal and Hermitian Green’s function. Here, $\hat{\pi}(x)=\frac{\partial\hat{\mathscr{L}}(x)}{\partial\dot{\hat{\phi}}(x)}=\dot{\hat{\phi}}(x)$ is the momentum operator conjugate to $\hat{\phi}$. We recall from Sec. III.1 that the field operator $\hat{\phi}(x)$ and its Hermitian conjugate $\hat{\phi}^{\dagger}(x)$ do not commute at equal times, i.e.,
$\hat{\phi}(t,\vec{x}),\hat{\phi}^{\dagger}(t,\vec{y})]\neq 0\>,$ (52)
since $\hat{\phi}^{\dagger}$ is not the field operator dual to $\hat{\phi}$. Indeed, it is the dual field operator $\hat{\tilde{\phi}}^{\dagger}$ that commutes with $\hat{\phi}$ at equal times:
$\hat{\phi}(t,\vec{x}),\hat{\tilde{\phi}}^{{\dagger}}(t,\vec{y})]=0\>,$ (53)
which is trivially consistent with Eq. (51b) by virtue of the pseudo-reality condition (37).
The causal Green’s functions solving (27) are the retarded and the advanced propagators:
$G_{R}(x,y)=\left\{\begin{array}[]{l}D(x,y)-D(y,x)\text{\>\>\>for\>\>\>}x^{0}>y^{0}\\ 0\text{\>\>\>for\>\>\>}y^{0}>x^{0}\end{array}\right$.\>, (54c)
$G_{A}(x,y)=\left\{\begin{array}[]{l}0\text{\>\>\>for\>\>\>}x^{0}>y^{0}\\ D(y,x)-D(x,y)\text{\>\>\>for\>\>\>}y^{0}>x^{0}\end{array}\right$.\>, (54f)
@@
$D(x,y)=\int\!\frac{{\rm d}^{3}\vec{p}}{(2\pi)^{3}}\>\frac{1}{2iE_{\vec{p}}}\,e^{-ip\cdot(x-y)}\>.$ (55)
@@
We will show below that the operator $\hat{\eta}_{P}$ can be written as a combination of an operator $\hat{\eta}_{g}$, which can be obtained as a perturbative expansion in the coupling $g$, and the ‘standard’ parity operator $\hat{\mathcal{P}}$. The ‘standard’ parity operation has the usual form
$\hat{\mathcal{P}}(t)\hat{\phi}(t,\vec{x})\hat{\mathcal{P}}^{-1}(t)=-\hat{\phi}(t,-\vec{x})\>,$ (56)
with
$\hat{\mathcal{P}}(t)=e^{\frac{\pi}{2}\int{\rm d}^{3}\vec{z}(\hat{\phi}(t,\vec{z})\hat{\pi}(t,-\vec{z})+\hat{\phi}(t,\vec{z})\hat{\pi}(t,\vec{z}))}\>.$ (57)
The operator $\hat{\eta}_{g}$ is less trivial, and the full calculation is provided in Appendix A. We are, however, looking for an expansion of the form
@@
$+ig^{2}\int^{t}_{t_{0}}{\rm d}y^{0}\int^{y^{0}}_{t_{0}}{\rm d}z^{0}\int{\rm d}^{3}\vec{y}\>{\rm d}^{3}\vec{z}\>G_{R}(x,y)G_{R}(y,z)G_{R}(y,z)\hat{\phi}(z)\>.$ (58)
Combining this with the ‘standard’ parity operator, gives us the required expansion in Eq. (47):
$\hat{\eta}_{g}(t,t_{0})\hat{\mathcal{P}}(t)\hat{\phi}(t,-\vec{x})$ $\hat{\mathcal{P}}^{-1}(t)\hat{\eta}_{g}^{-1}(t,t_{0})=-\hat{\eta}_{g}(t,t_{0})\hat{\phi}(t,\vec{x})\hat{\eta}_{g}^{-1}(t,t_{0})$
$=-\hat{\phi}(t,\vec{x})+ig\int^{t}_{t_{0}}{\rm d}y^{0}\int{\rm d}^{3}\vec{y}\>G_{R}(x,y)\hat{\phi}^{2}(y)$
$+g^{2}\int^{t}_{t_{0}}{\rm d}y^{0}\int^{y^{0}}_{t_{0}}{\rm d}z^{0}\int{\rm d}^{3}\vec{y}\>{\rm d}^{3}\vec{z}\>G_{R}(x,y)\hat{\phi}(y)G_{R}(y,z)\hat{\phi}^{2}(z)$
$-ig^{2}\int^{t}_{t_{0}}{\rm d}y^{0}\int^{y^{0}}_{t_{0}}{\rm d}z^{0}\int{\rm d}^{3}\vec{y}\>{\rm d}^{3}\vec{z}\>G_{R}(x,y)G_{R}(y,z)G_{R}(y,z)\hat{\phi}(z)\>.$ (59)
As we will now show, the last term is a loop correction of order $\hbar$ and thus vanishes in the semi-classical limit, $\hbar\mapsto 0$, relevant to Eq. (47).
Taking the expectation value of Eq. (III.6) and using Wick’s theorem, we obtain
$-\big<\hat{\eta}_{g}(t,t_{0})\hat{\phi}(t,\vec{x})$ $\hat{\eta}_{g}^{-1}(t,t_{0})\big>=-\phi(t,\vec{x})-ig\int^{t}_{t_{0}}{\rm d}y^{0}\int{\rm d}^{3}\vec{y}\>G_{R}(x,y)[\phi^{2}(y)+\hbar G_{F}(y,y)]$
$-g^{2}\int^{t}_{t_{0}}{\rm d}y^{0}\int^{y^{0}}_{t_{0}}{\rm d}z^{0}\int{\rm d}^{3}\vec{y}\>{\rm d}^{3}\vec{z}\>G_{R}(x,y)G_{R}(y,z)$
$\times[\phi(y)\phi^{2}(z)+\hbar\phi(y)G_{F}(z,z)+2\hbar G_{F}(y,z)\phi(z)]$
$+i\hbar g^{2}\int^{t}_{t_{0}}{\rm d}y^{0}\int^{y^{0}}_{t_{0}}{\rm d}z^{0}\int{\rm d}^{3}\vec{y}\>{\rm d}^{3}\vec{z}\>G_{R}(x,y)G_{R}(y,z)G_{R}(y,z)\phi(z)\>,$ (60)
where we have reinstated factors of $\hbar$ for clarity and
$G_{F}(x,y)$ $=\big<\mathrm{T}[\hat{\phi}(x)\hat{\phi}(y)]\big$>
$=i\theta(x^{0}-y^{0})D(x,y)+i\theta(y^{0}-x^{0})D(y,x)+\mathcal{O}(g)$
$=\int\frac{{\rm d}^{4}p}{(2\pi)^{4}}\frac{ie^{-ip\cdot(x-y)}}{p^{2}-m^{2}+i\epsilon}+\mathcal{O}(g)$ (61)
is the Feynman propagator. (Note that the Feynman propagator $G_{F}$ is the full, dressed Feynman propagator, whereas the retarded propagator $G_{R}$ is tree level.) It is now clear that the terms of order $\hbar$ in Eq. (III.6) can be identified as loop corrections. As we take the classical limit $\hbar\mapsto 0$, these terms vanish, and we recover the classical expansion in Eq. (47) after identifying the $c$-number field $\phi(x)$ with $\big<\hat{\phi}(x)\big$> in the same limit. Notice that the combination $iG_{R}(y,z)[2G_{F}(y,z)-iG_{R}(y,z)]=G_{F}^{2}(y,z)+D^{2}(y,z)$, obtained from terms in the second and third lines, is nothing other than the retarded bubble diagram, as we would expect. The non-commutativity of the field and its Hermitian conjugate is now readily confirmed, and, at order $g$, we have
$\left[\hat{\phi}(t,\vec{x}),\hat{\phi}^{{\dagger}}(t,\vec{y})\right]=-ig\int{\rm d}^{4}z\;G_{R}(x,z)G_{R}(y,z)\hat{\phi}(z)\;,$ (62)
where $x^{0}=y^{0}=t$.
Given that the perturbation series for the solution for the transformation [Eq. (47)] is similar to that of the interacting Klein-Gordon equation, we find (see Appendix A) that the series expansion is generated by an operator similar to the time-evolution operator; specifically,
$\hat{\eta}_{g}(t,t_{0})=\hat{U}^{-1}\quantity(-\frac{1}{3},t,t_{0})\>,$ (63)
@@
$\hat{U}\quantity(\alpha,t,t_{0})\coloneqq T\quantity[e^{\alpha g\int^{t}_{t_{0}}{\rm d}y^{0}\int{\rm d}^{3}\vec{y}\>\hat{\phi}^{3}(y^{0},\vec{y})}]\>,$ (64)
and
$\hat{U}^{-1}\quantity(\alpha,t,t_{0})=\bar{T}\quantity[e^{-\alpha g\int^{t}_{t_{0}}{\rm d}y^{0}\int{\rm d}^{3}\vec{y}\>\hat{\phi}^{3}(y^{0},\vec{y})}]\>,$ (65)
for some constant $\alpha\in\mathbb{C}$. Here, $T$ is the time-ordering operator and $\bar{T}$ is the anti-time-ordering operator. Note that $\hat{U}(\frac{1}{3!},t,t_{0})$ is (when replacing $\hat{\phi}$ by $\hat{\phi}_{0}$) the usual time-evolution operator whose action on the free Klein-Gordon field $\phi_{0}$ gives the series expansion in Eq. (26).
@@
$\hat{\mathcal{P}}^{\dagger}(t)=\hat{\eta}_{P}(t,t_{0})\hat{\mathcal{P}}(t)\hat{\eta}_{P}^{-1}(t,t_{0})\;.$ (67)
@@
and the operator $\hat{\eta}_{g}$ in Eq. (63). Thus, the Lagrangian is anti-pseudo-Hermitian with respect to the anti-linear operator $\hat{\eta}_{T}$:
$\hat{\eta}_{T}(t,t_{0})=\hat{\eta}_{g}(t,t_{0})\hat{\mathcal{T}}(t).$ (70)
@@
Given that the Lagrangian is $\eta_{P}^{-1}\eta_{T}$-symmetric [see Eq. (41)], we see that it is also $PT$ symmetric with respect to the ‘standard’ $PT$ transformation, since $\hat{\eta}_{P}^{-1}\hat{\eta}_{T}=\hat{\mathcal{P}}\hat{\mathcal{T}}$.
@@
Thus, we have demonstrated that by adopting a first-principles approach and introducing the pseudo-reality condition, we can achieve a self-consistent formulation of the $i\phi^{3}$ theory. This framework resolves the contradictions arising from Euler-Lagrange equations and ‘standard’ parity and time-reversal transformations and ensures that the theory remains well defined with a real spectrum under exact $PT$ symmetry.
IV Generalisation to higher spins